Grade 12th passAlgebra

a=x^1/3.y^1/3+x^(-1/3).y^(-1/3) find a^3-3ay-x=?(A) y/x(B)x/y(C)x^2/y(D)y^2/x

Kapil
9 Years agoGrade 12th pass

4 Answers

Prince Khan
9 Years ago
Simply take x = 8 and y = 1 you will get A=2.5Now a³-3ay-x = 15.625-7.5-8=.625Put the value in optionsY²/x = 1/8 = .625
km mukul
8 Years ago
Simply take x = 8 and y = 1 you will get A=2.5Now a³-3ay-x = 15.625-7.5-8=.125Put the value in optionsY²/x = 1/8 = .125
km mukul
8 Years ago
Simply take x = 27and y = 1 you will get A=10/7 Now a³-3ay-x = 1000/27-10-27=1/27 Put the value in optionsY²/x = 1/27
xxxxx
8 Years ago
so interestingly you all have given answers .lets just put x=1 and y=27 ,you will get value as -233.96296296 what should i opt now?